# Nighttime Person Re-Identification via Collaborative Enhancement Network with Multi-domain Learning

### v1

@@ Abstract
[…]
1 Introduction
@@
Person re-identification (ReID) is the task of identifying individuals in the images captured by different surveillance cameras, and it has significant potential applications in the fields of public security and criminal investigations. Despite current state-of-the-art methods [7, 34, 19, 35, 2, 31, 9, 33] achieve excellent performance in ReID during the daytime, the challenges of re-identifying person in nighttime scenarios remains unsolved. The main reason can be attributed to the severe image degradation and reduced visibility of nighttime person images. Hence, a straightforward approach to address the nighttime person ReID problem commonly involves the application of a relighting technique to enhance the visibility of input images before feeding them into the ReID network, as illustrated in Figure 1 (a). However, this method is not always effective since the conventional relighting technique is not specifically tailored to the ReID task.
@@
Figure 1: Comparison of three nighttime person ReID strategies. (a) uses Zero_DCE [8] for low-light image pre-processing and TransReID [9] for recognition. (b) represents the current state-of-the-art sequential end-to-end nighttime ReID method IDF [18]. (c) introduces our proposed parallel end-to-end method, which outperforms the other strategies. Evaluation is performed on the nighttime ReID dataset Night600 [18].
@@
To overcome this limitation, existing works [36, 18] propose end-to-end nighttime person ReID methods that integrate relighting and ReID within a single framework, as depicted in Figure 1 (b). For instance, Zhang et al. [36] propose an end-to-end network that optimizes both the denoising and ReID objectives, leading to effective performance improvement compared to baseline methods. Nevertheless, their approach overlooks the challenges posed by low-light conditions in nighttime images. Lu et al. [18] propose an illumination distillation network that combines features from nighttime images and relighting images. This method aims to mitigate the limitations of solely relying on relighting images. Although these methods demonstrate promising performance improvements, they still follow a sequential approach of integrating relighting and ReID networks. Consequently, the performance of the ReID network is not only constrained by the quality of the relighting image, but also lacks collaborative modeling between in two networks. Moreover, most existing relighting networks [8, 15, 37] simply stack multiple convolutional layers, making it challenging to capture the crucial global context necessary for modeling lighting variations.
@@
To address these challenges, we introduce the Collaborative Enhancement Network (CENet) that integrates ReID and relighting networks in parallel, as depicted in Figure 1 (c). This parallel structure design mitigates the impact of relighting image quality on the ReID model, and the designed multi-level feature interaction achieves effective collaborative modeling in ReID and re-illumination networks. CENet comprises three primary components: a shared encoder, a relighting subnet, and a ReID subnet. These components collaborate to learn lighting-enhanced features for both relighting and ReID tasks. Benefiting from relighting task-driven, the shared encoder can extract lighting-enhanced features from nighttime images, which helps to enhance the low-level ReID feature. Furthermore, feature distillation scheme distills lighting information from the decoder features of the relighting subnet to guide feature learning in the ReID subnet.
In the existing real nighttime dataset, we encounter challenges when training the proposed network due to its limited scale and the relighting network is hard to obtain corresponding daytime images for learning. To this end, we synthesize a large-scale nighttime dataset called Syn_Dark, which is created by introducing random factors, including illumination, color, and contrast variations, into two ReID datasets, Market1501 [42] and MSMT17 [28]. However, solely relying on synthesized data limits the network’s performance on real-world nighttime ReID datasets. To overcome this limitation, we design a multi-domain learning approach that combines both real and synthesized data domains. Our approach learning the network in the real domain using an unsupervised relighting method based on the RetineX theory. Simultaneously, in the synthetic domain data, which includes pairs of daytime ground-truth images, we employ a fully supervised method. During the iterative training process, we alternate between different domains and their respective learning approaches to comprehensively optimize the entire network. This strategy helps reduce the domain gap and improve the network’s performance on real-world nighttime ReID datasets.
In summary, the main contributions are as follows:
  • We propose a new parallel framework that addresses the impact of relighting image quality on the ReID task. Furthermore, our framework can get rid of relighting subnet during test time, avoiding additional inference costs.
  • We design an novel multilevel feature interaction strategy to facilitate collaborative modeling between the ReID and relighting networks, which enhances the ReID network’s feature representation capability specifically for nighttime images.
  • We present an efficient multi-domain learning strategy that effectively utilizes large-scale synthetic data to overcome the limitations of model training with small-scale real night data.
2 Related Work
2.1 Nighttime Person ReID Methods
@@
Due to the challenges posed by low-light conditions at night, visible light cameras often struggle to capture clear images during nighttime. Some studies [29, 20] propose to use visible light devices during the day and introduce infrared devices at night to capture person images, named cross-modality person ReID. Recently, research in this field has attracted increasing attention [16, 32, 25, 39, 30, 33], with the primary challenge is reducing the differences between various modalities. To address this, Li et al.[16] employ a self-supervised learning network by establishing an intermediate $X$ modality to mitigate the modalities gap. Zhang et al.[39] develop a network that improves the consistency of features across different channels. Ye et al.[30] propose a strategy involving random color channel swapping to create images that are not dependent on color within the same modality. Other studies [10, 11, 40] synthesize low-light datasets to explore challenges of low-light images, which ignores synthetic data hard to reflect the real challenges of person ReID at night. To this end, Zhang et al. [36] introduce a nighttime person ReID dataset focusing on the noise of nighttime images, and propose a denoising-then-match framework. Then, Lu et al. [18] propose a real nighttime ReID dataset that considers complex lighting conditions and develops an enhancement-then-match framework. However, these methods are still constrained by the quality of relighting images. This paper presents a parallel architecture integrating relighting and ReID networks, which effectively avoids direct dependence on relighting images, and designs a multi-level feature enhancement scheme between the two networks.
2.2 Relighting Methods
@@
The relighting task aims to recover invisible regions in low-light images. Some classic methods [5, 22, 21, 13, 14, 12] adjust image brightness by different theories and provides an important foundation for future research. With the rise of deep learning, a series of methods combine convolutional neural networks and RetineX theory [27, 38, 37] to achieve more accurate enhancement effects. Wei et al. [27] first used RetineX theory with a deep neural network, employing a two-step approach of image decomposition followed by enhancement. Zhang et al. [38] design a simple and effective network that maps the illumination and reflection maps to different subspaces, enabling separate correction of brightness and degradation. However, these methods heavily rely on paired data for training, limiting their applicability. To address this limitation, Guo et al. [8, 15] design a network to generate pixel-level curve parameter maps, which are then used to adjust low-light images. Zhang et al. [37] propose an unsupervised relighting network based on information entropy and RetineX theory, which only uses low-light images for learning and outputs light and reflection maps. However, pretrained relighting networks cannot guarantee the optimal performance for person ReID tasks at night.
3 Methodology
@@
In this section, we introduce the proposed collaborative enhancement network in detail, including network overview, ReID branch, relighting branch, multilevel feature interactions, and multi-domain joint learning scheme.
Figure 2: (a) The architecture of the collaborative enhancement network for nighttime ReID. (b) The detail of the transformer decoder, where $\bigoplus$ represents the feature addition.
3.1 Network Overview
@@
The proposed network framework, as illustrated in Figure 2, comprises three main components: the shared encoder, ReID subnet, and relighting subnet. By combining the shared encoder with the ReID subnet and relighting subnet, we establish the ReID branch and relighting branch, enabling a parallel architecture for both vision tasks. To elaborate, we first utilize the shared encoder to extract features from nighttime person images. These features are then directed to two distinct task subnets, each producing its respective output results. The design of the shared encoder allows the ReID task to benefit from the low-level features relevant to the relighting task. In addition, we introduce a feature distillation mechanism to enhance the high-level feature representation in the ReID subnet. In contrast to existing methods, our proposed network achieves effective collaborative modeling between image relighting and ReID tasks, while avoiding the impact of enhanced image quality on ReID performance.
3.2 ReID Branch
@@
Feature Extraction. The ReID branch adopts the Transformer network ViT [6] with strong representation ability as the feature extraction network. Within this branch, the initial 5 encoders are designated as the shared encoder, while the remaining 6 encoders and task head serve as the ReID subnet. We partition the nighttime image $X$ into $N$ equally-sized image patches denoted as ${X_{p}^{i}|i=1,2,...,n}$, incorporating a trainable global token $X_{cls}$ as the input sequence for the ReID branch. Additionally, drawing from the TransReID [9], we incorporate a learnable position embedding into the input sequence. Moreover, we introduce camera information embedding to enhance the acquisition of invariant features within the input sequences. Subsequently, the input sequence is passed to the shared encoder, yielding the shared features, which can be expressed as:
$F_{share}=\left[\mathcal{E}\left(x_{cls}\right);\mathcal{E}\left(x_{p}^{1}\right);\mathcal{E}\left(x_{p}^{2}\right);\cdots;\mathcal{E}\left(x_{p}^{N}\right)\right]$ (1)
@@
where $\mathcal{E}$ represents the shared encoder, and $F_{share}$ corresponds to the output features produced by the shared encoder. Given that ReID is a classification task, extracting deep high-level semantic features is essential for distinguishing target individuals. To achieve this, we establish a ReID subnet comprising 6 Transformer encoders to extract more discriminative identity features. Finally, we employ a fully connected layer to adjust the number of extracted identity features for the final classification.
@@
Task Learning. We employ triplet loss and identity loss for joint optimization. In specific, the identity loss is the cross-entropy loss without label smoothing, which can be formulated as:
$\mathcal{L}_{ID}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{N}y_{i,j}\ log{(\frac{e^{\alpha(s_{i,j}-m)}}{e^{\alpha(s_{i,j}-m)}+\sum_{k\neq j}e^{\alpha s_{i,k}}})}$ (2)
@@
where $B$ and $N$ represent the total number of samples and identities, respectively. Additionally, $y_{i,j}$ denotes the label assigned to the $j$-th identity of the $i$-th sample, while $s_{i,j}$ represents the score assigned to the $j$-th identity of the $i$-th sample. The variable $m$ represents the mean value of the identity score, and $\alpha$ is the temperature parameter. The triplet loss is defined as follows:
$L_{Tri}=\max(0,\alpha+d(f_{a},f_{p})-d(f_{a},f_{n}))$ (3)
@@
where $f_{a}$, $f_{p}$, and $f_{n}$ denote the feature representations of anchor samples, positive samples, and negative samples, respectively. Here, $d$ denotes the Euclidean distance, and $\alpha$ is a constant value used to control the feature margin. Furthermore, we optimize the ReID branch from both global and local perspectives, aiming to learn more fine-grained information. More detailed information in [9].
3.3 Relighting Branch

### v2

@@ Abstract
Prevalent nighttime person re-identification (ReID) methods typically combine image relighting and ReID networks in a sequential manner. However, their performance (recognition accuracy) is limited by the quality of relighting images and insufficient collaboration between image relighting and ReID tasks. To handle these problems, we propose a novel Collaborative Enhancement Network called CENet, which performs the multilevel feature interactions in a parallel framework, for nighttime person ReID. In particular, the designed parallel structure of CENet can not only avoid the impact of the quality of relighting images on ReID performance, but also allow us to mine the collaborative relations between image relighting and person ReID tasks. To this end, we integrate the multilevel feature interactions in CENet, where we first share the Transformer encoder to build the low-level feature interaction, and then perform the feature distillation that transfers the high-level features from image relighting to ReID, thereby alleviating the severe image degradation issue caused by the nighttime scenario while avoiding the impact of relighting images. In addition, the sizes of existing real-world nighttime person ReID datasets are limited, and large-scale synthetic ones exhibit substantial domain gaps with real-world data. To leverage both small-scale real-world and large-scale synthetic training data, we develop a multi-domain learning algorithm, which alternately utilizes both kinds of data to reduce the inter-domain difference in training procedure. Extensive experiments on two real nighttime datasets, Night600 and RGBNT201rgb, and a synthetic nighttime ReID dataset are conducted to validate the effectiveness of CENet. We release the code and synthetic dataset at: https://github.com/Alexadlu/CENet.
Index Terms:
Nighttime person ReID, parallel structure, illumination enhancement, feature distillation, multi-domain learning.
I Introduction
@@
Person re-identification (ReID) is the task of identifying individuals in the images captured by different surveillance cameras, and it has significant potential applications in the fields of public security and criminal investigation. Despite current state-of-the-art methods [1, 2, 3, 4, 5] achieve excellent performance during the daytime, the challenges of person ReID in nighttime scenario remain unsolved. Specifically, unlike existing ReID studies that mainly focus on challenges such as pedestrian viewpoint differences and pedestrian occlusion, nighttime person ReID additionally introduces some challenges such as low light, imaging noise, and color shift caused by the limitations of imaging devices in nighttime scenarios. It raises significant challenges for existing ReID algorithms. A straightforward approach to address the nighttime person ReID problem commonly involves the application of image relighting technique to enhance the visibility of input images before feeding them into ReID model, as shown in Figure 1 (a). However, these methods are not always effective because conventional relighting techniques are not specifically tailored to the ReID task.
@@
Figure 1: Comparison of three nighttime person ReID strategies. (a) uses Zero_DCE [6] for low-light image pre-processing and TransReID [4] for recognition. (b) represents the current state-of-the-art sequential end-to-end nighttime ReID method IDF [7] with. (c) introduces our proposed parallel end-to-end method, which outperforms the other strategies. (b) and (c) have consistent training strategy and training data to accurately reflect advantages of the parallel framework. The Rank values of the three strategies are evaluated on the same nighttime ReID dataset Night600 [7]
@@
To overcome this limitation, existing works [8, 7] propose end-to-end nighttime person ReID methods that integrate image relighting and ReID within a single framework, as depicted in Figure 1 (b). For instance, Zhang et al. [8] propose an end-to-end network that optimizes both the denoising and ReID objectives, leading to effective performance improvement compared to baseline method. Nevertheless, their approach overlooks the challenges posed by low-light conditions in nighttime images. Lu et al. [7] propose an illumination distillation network that combines features from nighttime images and relighting images to mitigate the limitations of solely relying on relighting images. Although these methods demonstrate promising performance improvements, they still follow a sequential framework of relighting and ReID networks. Consequently, the performance of ReID networks is not only constrained by the quality of relighting images, but also lacks effective collaborative modeling between these two tasks. Moreover, most existing relighting networks [6, 9, 10] simply stack multiple convolutional layers, making it challenging to capture the crucial global context to handle lighting variations.
@@
To address these challenges, we propose the Collaborative Enhancement Network (CENet) that integrates ReID and relighting networks in parallel, as depicted in Figure 1 (c). Such parallel structure design mitigates the impact of relighting image quality on ReID models, and also allows the multilevel feature interaction for effective collaborative modeling between ReID and relighting networks. The effective interaction between two heterogeneous tasks (relighting task and ReID task) is key to building collaborative enhancement. To this end, we propose a multilevel feature interaction scheme. In the low-level features, we adopt the current widely accepted shared-parameter design [11, 12] to achieve collaborative enhancement among low-level features. In high-level features, we innovatively design a feature distillation technique to achieve collaborative enhancement among high-level features while mitigating the impact on task-specific information. By incorporating feature interactions at both low and high levels, we fully exploit the collaboration between these two tasks. In particular, CENet comprises three primary components, including a shared encoder, a relighting subnet, and a ReID subnet. These components collaborate to learn lighting-enhanced features for both relighting and ReID tasks. Driven by relighting task, the shared encoder can extract lighting-enhanced features from nighttime images, which helps to enhance the low-level ReID features. In addition, to further utilize the lighting-enhanced features, a brightness feature distillation scheme is introduced from the relighting subnet to the ReID subnet to improve the high-level feature representations of ReID task.
Constrained by the annotation difficulty of nighttime images, existing nighttime ReID datasets are small scale, which is hard to train big networks such as Transformer. In this work, we synthesize a large-scale nighttime dataset called Syn_Dark, which is created by introducing some random factors including illumination, color, and contrast variations into two ReID datasets, Market1501 [2] and MSMT17 [3]. Although introducing large-scale synthetic datasets is a straightforward idea, the gap between real and synthetic data makes the learning of large-scale synthetic data inadequate. To handle this issue, we propose a novel multi-domain learning algorithm which designs different learning strategies tailored to the characteristics of real and synthetic domains, respectively. It can effectively exploit large-scale synthetic data to improve the performance of nighttime person ReID task. In particular, we design tailored learning schemes for different domains, enabling us to leverage data from two distinct domains concurrently within a unified network. Through alternately optimizing the network parameters in different domain data, the network can achieve effective learning of domain-invariant features, thus mitigating inter-domain gaps. The feasibility of this scheme is validated in several domain adaptation studies [13, 14, 15]. In addition, we introduce the multi-dimensional degradation strategy in synthetic data construction to improve the diversity of low-light scenes. We will release the synthetic dataset to public for free academic usage.
In summary, the main contributions are as follows.
  • We propose a novel parallel framework to alleviate the negative impact and enhance the positive impact of the image relighting task on the person ReID task. We release the code and synthetic dataset at: https://github.com/mmic-lcl.
  • We design a new multilevel feature interaction scheme to facilitate collaborative modeling between the ReID and relighting networks. The shared low-level encoder aims to extract the collaborative features and the feature distillation in high-level features enhances the semantic information from relighting network to ReID network.
  • We develop an effective multi-domain learning algorithm to utilize both advantages of large-scale synthetic and small-scale real data. The network optimization is alternately performed between synthetic and real domains, and their domain gap is thus alleviated in the training procedure.
The rest of this paper is organized as follows: Section II reviews related works. Section III introduces the methodology of CENet, including its architecture, multilevel feature interaction, and multi-domain learning. Section IV presents experimental results and comparisons. Finally, Section V concludes the paper.
II Related Work
II-A Person ReID Methods
Due to challenges posed by low-light conditions, visible light cameras often struggle to capture clear images at night. Some studies [16, 17] propose to use visible light devices during the day and introduce infrared devices to capture person images, named cross-modality person ReID. However, cross-modality person ReID arises one of the major challenges is the modality gap between visible and infrared images. Therefore, many efforts [18, 19, 20, 21, 22, 23, 24, 1, 25, 26, 27, 5] attempt to handle this issue. For example, Wang et al.[18] introduce a novel dual-level discrepancy reduction learning scheme, which handles the cross-modality discrepancies in image-level and feature-level, respectively. Choi et al.[19] propose a hierarchical cross-modality disentanglement method to automatically disentangle ID-discriminative factors and ID-excluded factors from visible-thermal images. Li et al.[20] design an auxiliary X modality and reformulate infrared-visible dual-modality cross-modality learning as an X-Infrared-Visible three-modality learning problem. Zhang et al.[21] design a diverse embedding expansion network to effectively generate diverse embedding for nighttime cross-modality person ReID. Cui et al.[28] adopt a dual-modality transfer module to perform compensation for information asymmetry in HSV color space, which can effectively alleviate cross-modality discrepancies and better preserve discriminative identity features.
Li et al.[22] employ a self-supervised learning network by establishing an intermediate modality to mitigate the modalities gap. Similarly, Ye et al.[23] reduce modality differences through an intermediate modality and employ a tri-modality joint learning strategy. To address the issue of potentially noisy labels in nighttime data annotations, Ye et al.[5] mine the cues at the channel-level, part-level, and graph-level. Zhang et al.[1] develop a network to improve the consistency of features across different channels. Ye et al.[25] propose a strategy involving random color channel swapping to create images that are not dependent on color within the same modality.
@@ II-A Person ReID Methods
Other studies [29, 30, 31] synthesize low-light datasets to explore challenges of low-light images, but synthetic data are hard to reflect the real challenges of nighttime person ReID. To this end, Zhang et al. [8] introduce a nighttime person ReID dataset focusing on the noise of nighttime images, and propose a denoising-then-match framework. Then, Lu et al. [7] propose a real nighttime ReID dataset that considers complex lighting conditions and develops an enhancement-then-match framework. However, these methods are still constrained by the quality of relighting images. In this work, we present a parallel architecture integrating relighting and ReID networks, which effectively avoids direct dependence on relighting images and also facilitates introducing the multilevel feature enhancement scheme between the two networks.
Moreover, the color shift issue, caused by imaging bias in clothes color under different low-light conditions, is also a challenge in nighttime person ReID. The challenge is actually similar to a challenge in clothing-change ReID. Existing studies on clothing-change ReID already achieve relatively mature progress. For instance, Han et al. [32] propose a clothing-change feature augmentation model that significantly expands the clothing-change data in feature space rather than visual image space. Yang et al. [33] present an auxiliary-free competitive identification model, which achieves a win-win situation by enriching the identity-preserving information conveyed by the appearance and structure features while maintaining holistic efficiency. This design might also benefit for nighttime person ReID. It is worth mentioning that nighttime data are difficult to annotate, and thus acquiring large-scale annotated nighttime person ReID data is extremely difficult. However, existing studies on unsupervised ReID provide some reference value for solving the learning problem of unlabelled data at night. For instance, Chen et al. [34] propose a dual clustering co-teaching approach, which exploits a two-stream network to extract features and a dual clustering strategy to generate two sets of pseudo-labels for network training.
II-B Image Relighting Methods
The image relighting task aims to recover invisible regions in low-light images. A classic method for adjusting image brightness is Histogram Equalization (HE) [35], which globally modifies the pixel distribution. However, it overlooks local information, potentially leading to loss of details, amplified noise, and suboptimal color reproduction. Some classic methods [35, 36, 37, 38, 39, 40] adjust image brightness by different theories and provide an important foundation for future research. Hence, various enhanced techniques [36, 37, 38] are proposed to further improve enhancement results. Additionally, the RetineX model [39], inspired by the human visual system, has gained widespread usage in enhancing low-light images. This model assumes that an image can be decomposed into an illumination map and a reflection map, with the reflection map, which represents the enhancement result.
Figure 2: The architecture of the Collaborative Enhancement Network (CENet) for nighttime person ReID. CENet consists of three parts: the shared encoder, the ReID subnet and the Relighting subnet. In addition, the dashed part will be removed during the inference phase.
@@ II-B Image Relighting Methods
With the rise of deep learning, a series of methods combine convolutional neural networks and RetineX theory [41, 42, 10, 43, 44] to achieve more accurate enhancement effects. Wei et al. [41] first use RetineX theory with a deep neural network, employing a two-step approach of image decomposition followed by enhancement. Zhang et al. [42] design a simple and effective network that maps the illumination and reflection maps to different subspaces, enabling separate corrections of brightness and degradation. However, these methods heavily rely on paired data for training, limiting their applicability. To address this limitation, Guo et al. [6, 9] design a network to generate pixel-level curve parameter maps, which are then used to adjust low-light images. Zhang et al. [10] propose an unsupervised relighting network based on information entropy and RetineX theory, which only uses low-light images for learning and outputs light and reflection maps. However, pretrained relighting networks cannot guarantee the optimal solution for nighttime person ReID.
In addition, we further discuss several methods that can perform global lighting in computer vision. Specifically, existing methods according to different learning strategies can be divided into two categories: supervised learning-based and unsupervised learning-based methods. Among them, supervised learning-based relighting methods [45, 46] are current mainstream, but rely on large-scale paired data for training. However, it is extremely difficult to obtain paired data in real-world scenarios. Therefore, some studies [47] use synthetic techniques to build paired data for training. However, they do not yield satisfactory results due to the domain gap between real and synthetic data. In order to avoid paired data limitations, some researchers propose unsupervised relighting methods [48, 44], and usually design various reference-free losses to guide model optimization. However, these methods focus on improving the quality of visual perception, and thus using them as a pre-processing step does not always guarantee the performance of nighttime person ReID.
III Methodology
@@
In this section, we introduce the proposed Collaborative Enhancement Network (CENet) in detail, including network overview, ReID branch, relighting branch, multilevel feature interactions, and multi-domain learning scheme.
III-A Architecture Overview
@@
The proposed framework, as shown in Figure 2, comprises three main components, including the shared encoder, ReID subnet, and relighting subnet. We first utilize the shared encoder to extract features from nighttime person images. These features are then directed to two separate subnets of person ReID and image relighting. The design of the shared encoder allows the ReID task to benefit from the low-level features relevant to the relighting task. Then, we introduce a feature distillation mechanism to enhance the high-level feature representation of ReID task. In contrast to existing methods, our proposed network achieves effective collaborative modeling between image relighting and ReID tasks, while avoiding the impact of image lighting quality on ReID performance.
III-B ReID Branch
@@
The ReID branch adopts the Transformer network ViT [49] with strong representation ability as the backbone network. Within this branch, the initial five encoders are designated as the shared encoder, while the remaining six encoders and task head serve as the ReID subnet. We partition the nighttime image $X$ into $N$ equally-sized image patches denoted as ${X_{p}^{i}|i=1,2,...,n}$, incorporating a trainable global token $X_{cls}$ as the input sequence for the ReID branch. Additionally, drawing from the TransReID [4], we incorporate a learnable position embedding into the input sequence. Moreover, we introduce camera information embedding to enhance the acquisition of invariant features within the input sequences. Subsequently, the input sequence is passed to the shared encoder, yielding the shared features, which can be expressed as:
$F_{share}=\left[\mathcal{E}\left(x_{cls}\right);\mathcal{E}\left(x_{p}^{1}\right);\mathcal{E}\left(x_{p}^{2}\right);\cdots;\mathcal{E}\left(x_{p}^{N}\right)\right],$ (1)
@@
where $\mathcal{E}$ represents the shared encoder, and $F_{share}$ corresponds to the output features produced by the shared encoder. Given that ReID is a classification task, extracting deep high-level semantic features is essential for distinguishing target individuals. To this end, we establish a ReID subnet comprising 6 Transformer encoders to extract more discriminative identity features. Finally, we employ a fully connected layer to adjust the number of extracted identity features for the final classification.
@@
We employ the triplet loss and identity loss for joint optimization. In specific, the identity loss is the cross-entropy loss without label smoothing, which can be formulated as:
$\mathcal{L}_{ID}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{N}y_{i,j}\ log{(\frac{e^{\alpha(s_{i,j}-m)}}{e^{\alpha(s_{i,j}-m)}+\sum_{k\neq j}e^{\alpha s_{i,k}}})},$ (2)
@@
where $B$ and $N$ represent the total numbers of samples and identities, respectively. Additionally, $y_{i,j}$ denotes the label assigned to the $j$-th identity of the $i$-th sample, while $s_{i,j}$ represents the score assigned to the $j$-th identity of the $i$-th sample. The variable $m$ represents the mean value of the identity score, and $\alpha$ is the temperature parameter.
The triplet loss is defined as follows:
$\mathcal{L}_{Tri}=\max(0,\alpha+d(f_{a},f_{p})-d(f_{a},f_{n})),$ (3)
@@
where $f_{a}$, $f_{p}$, and $f_{n}$ denote the feature representations of anchor samples, positive samples, and negative samples, respectively. Here, $d$ denotes the Euclidean distance, and $\alpha$ is a constant value used to control the feature margin. Furthermore, we optimize the ReID branch from both global and local perspectives, aiming to learn more fine-grained information. More detailed information can be referred to [4].
III-C Relighting Branch